OpenStudy (anonymous):

lim x approaches pi/2 sin3x/x

OpenStudy (anonymous):

The function is continuous at \( x = \frac \pi 2\)

OpenStudy (anonymous):

Replace x by its value and you are done. \[ \sin\left( \frac {3 \pi}2 \right)=? \]

OpenStudy (anonymous):

still cannot understand answer. sinx/x =1 how to find sin 3x/x using pi/2 ?

OpenStudy (anonymous):

Limit of sinx/x =1 when x approaches zero which makes 0/0 consideration. Here we do not have this problem. In fact, it is given to trick you into using sinx/x, but you should not

OpenStudy (anonymous):

\[ \sin\left( \frac {3 \pi}2 \right)=-1\\ \]

OpenStudy (anonymous):

The limit is \[ \Large \frac {-1} {\frac {\pi}2}= - \frac 2 { \pi} \]